Two trains leave the depot at the same time, one traveling east and the other west. The speed of one train is 9mph slower than the other. If after 4 hours the distance between the trains is 196 miles find the speed of each train.: Two trains leave the depot at the same time, one traveling east and the other west. The speed of one train is 9mph slower than the other. If after 4 hours the distance between the trains is 196 miles find the speed of each train.
Answer by mangopeeler07(423)

x=speed of one train
x-9=speed of the other train
mph(hr)=total distance

4(x)+4(x-9)=196---------------------------------because the sum of 4 hours times both speeds will give you the total distance

Simplify
4x+4x-36=196

Combine like terms
8x-36=196

Add 36 to each side
8x=232

Divide by 8
x=29

[x=speed of one train]
[x-9=speed of the other train]

speed of one train= 29mph
speed of the other train= 20mph